A number is decreased by 30% to 154. What is the original amount?
Answer by jim_thompson5910(35256)

let x = unknown number

A number is decreased by 30% to 154

x is decreased by 30% to 154

x - 30% of x = 154

x - 0.3x = 154

0.7x = 154

x = 154/0.7

x = 220

So the original amount is 220